# Remote Sensing Image Intelligent Interpretation with the Language-Centered Perspective: Principles, Methods and Challenges

### v1

@@ III-B Language-Centered Unified Multimodal Representation
In language-centered remote sensing intelligent interpretation systems, the key to unified multimodal representation lies in the effective mapping of heterogeneous information from different modalities into the language space. This requires not only the unification of data representation formats across modalities but also the accurate alignment of the semantics conveyed by each modality.
@@
The structure of this section is illustrated in Fig.6, and representative methods corresponding to each component are summarized in Table I.
@@
TABLE I: Unified multimodal representation structure and representative methods.
@@
Dynamic Hybird Embedding Oscar[33], VIMA[34], PaLM-E[35], TEOChat[36], SkySenseGPT[37]
@@ III-B1 Structural Alignment Across Modalities
Data from different modalities often exhibit significant differences in spatial resolution and temporal coverage. For example, the spatial resolution of remote sensing imagery may vary depending on the sensor type (e.g., RGB, MSI, SAR), while temporal coverage is influenced by the frequency of data acquisition. Traditional approaches typically represent each modality in distinct formats, making cross-modal fusion difficult and lacking effective channels for information interaction. However, with the success of Transformers[69] in natural language processing, an increasing number of studies have demonstrated that it is possible to achieve unified representation of multimodal data by constructing modality-specific embeddings. In this approach, the Transformer processes data from different modalities at the embedding layer, effectively mapping them—despite spatial and temporal differences—into a shared semantic space, thus enabling seamless fusion and collaboration across modalities. Depending on how the embeddings from different modalities are combined, this strategy can be categorized into two main types: Static Sequential Embedding, and Dynamic Hybrid Embedding.
@@
As illustrated in Fig.7(a), the static sequential embedding approach combines embeddings from different modalities in a fixed order and length after passing each modality through its respective encoder. For example, in VisualBERT[28], the first N embeddings represent textual information, while the following K embeddings represent visual information, separated by a [SEP] token. This allows the model to process and interpret the fused multimodal data jointly. Building upon this, LXMERT[29] enhances the model’s capacity to extract both intra-modal and inter-modal features by incorporating self-attention and cross-attention mechanisms. UNITER[30] further improves multimodal interaction learning by combining four training objectives: Masked Language Modeling (MLM), Masked Region Modeling (MRM), Word-Region Alignment (WRA), and Image-Text Matching (ITM).
@@
In the field of remote sensing, RSGPT[31] integrates visual and language modalities using a Q-Former to fuse modality-specific embeddings. It introduces image-question-answer triplets based on remote sensing imagery interpretation as domain-specific knowledge, and proposes RSIEval as a benchmark for evaluating the performance of remote sensing multimodal large language models (RSMLLMs). Building on this, GeoChat[32] fine-tunes a pretrained LLaMA[16] model using four types of visual-language understanding tasks. It distinguishes visual and language modalities using <im_start> and <im_end> tokens, significantly improving the model’s understanding of remote sensing imagery. Similarly, RS-LLaVA[9] adopts the multimodal architecture of LLaVA[58], introducing a projection layer to map visual modalities. Remote sensing imagery and their descriptions are encoded as image tokens and text tokens, respectively, and the model is further optimized using LoRA[70] fine-tuning techniques.
@@
As shown in Fig.7(b), dynamic hybrid embedding refers to the approach in which data from different modalities are concatenated and directly input into the embedding layer for joint processing. Unlike the modality-tagging method, in dynamic hybrid embedding, the number and positions of embeddings from each modality are dynamically variable. This allows the model to effectively capture interaction patterns across modalities, thereby enhancing its understanding of multimodal joint representations. For example, OSCAR[33] directly combines textual data, object labels, and object images into a single input, which is uniformly processed through the embedding layer. This approach significantly improves the model’s ability to extract target information in query-based tasks. PaLM-E[35] builds on this by integrating continuous data streams from the physical world, enabling the model to interact with real-world environments. VIMA[34], focused on the domain of embodied intelligence, directly combines text tokens, object tokens, and action tokens, allowing the model to learn how to generate appropriate action feedback in response to different instructions and goals—thus enabling robots to perform a wide variety of tasks.
@@
In the field of remote sensing, TEOChat[36] employs a Shared Vision-Language Connector to achieve hybrid embedding of image tokens and text tokens, further enhancing the fusion of multimodal information. To improve the model’s understanding and reasoning ability with respect to geographic knowledge, SkySenseGPT[37] incorporates more detailed instructions into the text tokens and introduces both region-level and image-level scene graph information. This significantly strengthens the model’s ability to understand both local and global relationships among geo-objects. Through this approach, remote sensing models are better equipped to capture spatial information and object relationships across different scales and levels, thereby improving their overall capability for intelligent interpretation of remote sensing imagery.
@@
The shift from modality-tagging methods to joint embedding strategies has become a major development trend for improving the performance of language-centered multimodal models. The core idea is to increase the model’s learning difficulty by disrupting the order of tokens from different modalities, thereby encouraging deeper understanding of multimodal data. Modality-tagging methods assign fixed positions and lengths to different modalities, allowing the model to clearly distinguish the source of each token. However, this structured design can limit the model’s ability to capture complex relationships across modalities. In contrast, joint embedding mixes and concatenates tokens from different modalities in a less constrained manner, compelling the model to learn inter-modal interactions and complementary information within a more flexible and intricate input structure. As a result, the model must not only understand the individual characteristics of each modality but also learn how to integrate latent relationships across modalities—thus enhancing its overall capability to interpret multimodal data holistically.
@@ III-B2 Semantic Alignment Across Modalities
Aligning representations from different modalities at the semantic level is a fundamental challenge in multimodal alignment research. Existing approaches to semantic alignment can be broadly categorized into two types: explicit alignment and implicit alignment. The key difference between them lies in whether the alignment relationships across modalities are explicitly introduced during training—such as through additional manual annotations or pre-defined alignment schemes.
@@
To effectively embed different modalities into the language modality, a straightforward approach is to use paired data—such as image-text or point cloud-text datasets—to train the model. Within this training framework, the model learns to align modalities by modeling the pairing relationships between multimodal data and corresponding textual descriptions. Based on the granularity of alignment between non-linguistic and linguistic modalities, current research efforts can be broadly classified into three main approaches: Scene-Caption alignment, Region-Phrase alignment, and Scene-Instruction alignment. The characteristics of each approach are illustrated in Fig.8.
@@
  • Scene-Caption Alignment: Scene-Caption alignment methods focus on aligning an entire image with its corresponding textual description, aiming to help the model learn the overall association between visual content and language, as illustrated in Fig.8(a). A representative work in this category is CLIP[7], which leverages contrastive learning to achieve effective alignment and has inspired a series of follow-up methods, including UniCL[38], K-LITE[39], and LaCLIP[40]. Due to the task-agnostic nature of contrastive learning, these approaches are well-suited as foundational models that can be quickly adapted to various downstream tasks via fine-tuning, few-shot, or even zero-shot learning. However, these methods face several limitations. First, they rely heavily on large-scale, high-quality image-text paired datasets, and their image understanding capability tends to be shallow. Second, such methods often struggle to handle both image understanding and text generation tasks simultaneously. To address these issues, SimVLM[41] introduced the PrefixLM pretraining objective, which maintains strong image understanding while enabling text generation capabilities. Similarly, BLIP[20] performs well in both image understanding and captioning, and its unique CapFilt mechanism effectively reduces noise in the training data. Despite these improvements, Scene-Caption-based methods generally lack the ability to deeply and finely understand image content. This limitation largely stems from the fact that simple image-text pairs fail to provide sufficiently rich guidance, thereby constraining the model’s capacity to learn more complex vision-language relationships.
@@
    In remote sensing, vision-language models have rapidly become a research focus due to their superior capabilities in image understanding and language processing. One of the early efforts in Scene-Caption alignment for remote sensing was conducted by Qu et al.[42], who utilized the UCM Captions dataset. However, this dataset was relatively small in scale and contained simple descriptions, limiting its ability to represent complex content and diverse scenarios in remote sensing imagery. To address these limitations, subsequent works such as RSICD[43] and RSITMD[44] constructed larger, more diverse, and semantically richer datasets, while also developing task-specific models for various downstream applications. Building upon these foundations, RemoteCLIP[45] and GeoCLIP[46] have further advanced the field. These studies introduced large-scale Scene-Caption datasets for remote sensing containing up to 1.6 million and 5 million images, respectively, and proposed two foundational models tailored specifically for the remote sensing domain. Compared to earlier efforts, these studies adopt Transformer-based architectures, demonstrating a deeper understanding of remote sensing image content. As a result, they have established new directions and benchmarks for vision-language model research in remote sensing.
@@
  • Region-Phrase Alignment: To enhance the model’s fine-grained understanding of images, some studies have explored incorporating detailed correspondences between image regions and individual words or phrases, as illustrated in Fig.8(b). Early work in this research direction—such as ViLBERT[47] typically relied on external object detectors to identify regions of interest (ROIs), using techniques like RoIAlign to extract features from these regions and align them with corresponding text tokens. However, this approach has several notable limitations. First, its dependence on pre-trained external detectors restricts the range of detectable object categories and directly impacts model performance. Second, the inclusion of an additional detection module increases model complexity and reduces training efficiency. To address these issues, VL-BERT[49] integrates the detection module into the training process, while ViLT[50] eliminates it altogether, instead using a pure Transformer architecture to extract image features. These adjustments significantly improve efficiency while maintaining strong performance. Additionally, some researchers argue that relying solely on ROI features may lead to the loss of valuable contextual information in the background. In response, strategies such as background blurring, masking, and feature fusion have been proposed to preserve as much background information as possible while extracting regional features. These strategies further enrich the model’s understanding of image content, especially when dealing with complex scenes containing rich background and multi-layered information.
@@
    In remote sensing, Region-Phrase alignment has emerged as a compelling research direction, particularly within the context of Scene-Caption tasks. Early studies primarily adopted traditional encoder-decoder architectures, focusing on extracting features from target regions for image caption generation. While these approaches proved effective, they often overlooked domain-specific knowledge unique to remote sensing. With the rapid advancement of vision-language models, concept-level Region-Phrase alignment has seen new breakthroughs in remote sensing. For instance, GeoVG[55] was the first to introduce the Visual Grounding task into the remote sensing domain. It employs a dual-stream architecture to separately extract visual features and geographic relational graphs, which are then integrated through a fusion module to effectively combine geospatial relationships with visual information. Building upon this, RSVG[56] introduces the MGVLF module, which enhances grounding accuracy by filtering out noise and reinforcing salient features. Additionally, Few-Shot Object Detection (FSOD) has become an emerging topic in remote sensing. A recent work, TEMO[57], has made innovative contributions in this direction by incorporating detailed descriptions of all object classes as additional features, effectively mitigating class confusion. These advancements not only demonstrate the significant potential of Region-Phrase alignment in the remote sensing domain, but also provide valuable insights and directions for future research.
@@
  • Scene-Instruction Alignment: While the previously discussed Scene-Caption and Region-Phrase alignment methods effectively achieve direct alignment between visual and language modalities, they often fall short in exploring the complex and abstract relationships among concepts expressed across modalities—an aspect that is crucial for achieving higher-level compositional generalization. Against this backdrop, Scene-Instruction-based alignment demonstrates its unique value. Its core lies in answering questions through logical reasoning, as illustrated in Fig.8(c). This requires the model not only to understand individual concepts but also to grasp the intricate relationships between them. In recent years, LLMs have demonstrated powerful general knowledge understanding and reasoning capabilities, which has driven a new research direction that leverages LLMs to tackle multimodal tasks. By fine-tuning LLMs with Scene-Instruction datasets—as done in LLaVA[58] and InstructBLIP[59] models can further enhance their logical reasoning and instruction-following abilities. These works have also simplified the process of constructing high-quality instruction datasets. Meanwhile, recent studies have introduced novel instruction processing mechanisms: for example, CAT[60] expands narrow language instructions into vision-language instructions, while SEAL[61] incorporates a visual search mechanism to help the model focus more precisely on key regions within an image. These innovations not only improve the model’s ability to understand complex scenes but also contribute new technical pathways toward the development of general artificial intelligence.
@@
    In remote sensing, Scene-Instruction alignment is gradually emerging as a critical technique. It enables non-expert users to interact with remote sensing imagery using natural language. In RSVQA[62], researchers used OpenStreetMap data and manual annotation to create the first large-scale remote sensing visual question answering (VQA) dataset, and proposed a CNN-RNN-based VQA model. Subsequent work advanced the field by developing VQA-TextRS[63], a dataset supporting open-ended question answering, alongside a Transformer-based VQA model. In the same year, Prompt-RSVQA[64] proposed a method for mapping images into the language space and feeding them, along with questions, into BERT[71] to generate answers. These studies indicate that remote sensing research is rapidly converging toward the direction of vision-language models, laying the groundwork for future innovations in human-AI interaction and multimodal scene understanding.
@@
In summary, the progression from Scene-Caption to Region-Phrase, and more recently to Scene-Instruction alignment reflects a gradual deepening in the semantic understanding of geo-objects within remote sensing intelligent interpretation. This evolution corresponds to a shift from coarse- to fine-grained image-text matching strategies. Specifically: Scene-Caption alignment focuses on matching entire images with holistic textual descriptions, representing a coarse-to-coarse granularity alignment. While this approach captures the overall semantics of an image, remote sensing imagery often contains multiple geo-objects, and global descriptions may struggle to adequately and precisely express the visual content present in the image. Region-Phrase alignment refines the image side by dividing it into multiple regions and aligning each with a corresponding phrase. Here, the image is fine-grained while the text remains relatively coarse (describing entire regions). This method allows for more accurate localization and description of specific geo-objects in the image, enabling multi-object interpretation. However, it lacks detailed semantic modeling of conceptual relationships between geo-objects. Scene-Instruction alignment further enhances the granularity of image-text alignment by using complex instructions to describe the relationships among geo-objects and guide the model’s attention across multiple objects in the image. This approach not only aligns multiple targets in the imagery but also encodes rich semantic instructions that express inter-object relationships, thereby deepening the model’s interpretive understanding of remote sensing scenes. The development of image-text alignment methods—from coarse to fine granularity—has accelerated the evolution of remote sensing interpretation models from shallow semantic understanding toward deep mechanistic comprehension.
@@
While explicit alignment methods have achieved remarkable success in multimodal learning, they also present several practical limitations. First, they rely on large-scale paired image-text datasets for model training, which poses significant challenges in terms of data collection and preprocessing. Second, achieving alignment between two different modalities typically requires separate training procedures for each modality, increasing both computational and time costs. More importantly, as the number of modalities increases, even when using a language-centered alignment strategy, the alignment process becomes exceedingly complex and cumbersome. Moreover, when interaction is required between two non-language modalities, using language as an intermediary can lead to issues in both efficiency and accuracy. If the alignment between one modality and language is imperfect, errors may accumulate during the conceptual transfer process, ultimately leading to suboptimal or incorrect interaction outcomes.
@@
Despite the significant heterogeneity across different modalities, they ultimately describe the same physical world. This fundamental fact implies that, regardless of whether the alignment between visual and language modalities is explicitly defined, there should exist an intrinsic condition for unified convergence. In support of this, Minyoung Huh et al.[72] experimentally demonstrated that visual models with different architectures and training objectives exhibit a high degree of representational consistency, and that this consistency continues to improve as model size and training data scale increase. This pattern also holds for cross-modal vision-language models. Similarly, Maniparambil[65] showed that it is possible to perform subtitle matching or retrieval tasks using unaligned visual and language encoders—further validating the existence of an inherent consistency between visual and language representation spaces. These findings raise a fundamental question: Is there an implicit, inherently defined alignment between language and vision? If so, then it may be possible to achieve concept-level alignment between vision and language even when training on single-modal data—this is the central idea behind implicit alignment, which has recently garnered increasing research interest. Notably, many of these studies choose to align visual representations to the language space, rather than the other way around. This leads to a second key question: Why is language often chosen as the convergence point for vision-language alignment? This paper argues that, as the primary medium of human communication, language most accurately reflects human information cognition and logical reasoning mechanisms. Furthermore, language is inherently abstract, generalizable, and adaptable, making it well-suited as the global workspace for efficiently organizing a conceptual space that encompasses visual and other non-linguistic modalities. Building on this view, this paper refers to such research efforts as language-centered implicit multimodal alignment.
@@
Language-centered implicit multimodal alignment aims to explore how non-linguistic modality information can be directly aligned with the language representation space in the absence of prior modality-language alignment pairs. In this line of research, the work LM4VisualEncoding[66] discovered that even when trained solely on textual data, the encoder of a large language model demonstrates surprisingly strong performance on both pure visual tasks and multimodal tasks. This finding offers new insights into the use of language models for visual tasks, highlighting their ability to encode and process visual-language information through shared parameters. Furthermore, the recent LAMP[67] study confirms the potential of pretrained language models in handling 3D point cloud data, providing additional evidence of the broad adaptability of language models in interpreting and processing non-traditional textual data. Collectively, these research efforts have pushed the boundaries of applying large language models to computer vision tasks and expanded our understanding of their latent capabilities in processing unimodal visual data.
@@
With the rapid advancement of remote sensing and related technologies, vast amounts of remote sensing data are generated daily. However, constructing large-scale image-text paired datasets in the remote sensing domain remains a significant challenge. Therefore, introducing implicit alignment into remote sensing is both necessary and timely. Recently, several studies have begun to explore this direction. For example, GRAFT[68] leverages internet images captured from the same ground locations as remote sensing images to serve as an intermediary between remote sensing imagery and language. It proposes a method to train a remote sensing vision-language model without relying on any textual annotations. Such attempts to introduce implicit alignment into the remote sensing field provide promising new directions for addressing the shortage of large-scale multimodal image-text pairs. Nonetheless, this line of research is still in its early stages and requires further exploration in the future.
III-C Language-Centered Knowledge Association

### v2

@@ III-B Language-Centered Unified Multimodal Representation
In language-centered remote sensing intelligent interpretation systems, the key to unified multimodal representation lies in the effective mapping of heterogeneous information from different modalities into the language space. This task requires not only the unification of data representation formats across modalities but also the accurate alignment of the semantics conveyed by each modality.
@@
The structure of this section is illustrated in Fig.6, and representative methods corresponding to each component are summarized in Table II.
@@
TABLE II: Unified multimodal representation structure and representative methods.
@@
Dynamic Hybrid Embedding Oscar[87], VIMA[64], PaLM-E[35], TEOChat[59], SkySenseGPT[108]
@@ III-B1 Structural Alignment across Modalities
Data from different modalities often exhibit significant differences in spatial resolution and temporal coverage. For example, the spatial resolution of remote sensing imagery may vary depending on the sensor type (e.g., RGB, MSI, or SAR), while temporal coverage is influenced by the frequency of data acquisition. Traditional approaches typically represent each modality in distinct formats, making cross-modal fusion difficult and lacking effective channels for information interaction. However, with the success of transformers[163] in natural language processing, an increasing number of studies have demonstrated that it is possible to achieve unified representation of multimodal data by constructing modality-specific embeddings. In this approach, the transformer processes data from different modalities at the embedding layer, effectively mapping them—despite spatial and temporal differences—into a shared semantic space, thus enabling seamless fusion and collaboration across modalities. This strategy can be categorized into two main types depending on how the embeddings from different modalities are combined: static sequential embedding and dynamic hybrid embedding.
@@
As illustrated in Fig.7(a), the static sequential embedding approach combines embeddings from different modalities in a fixed order and length after passing each modality through its respective encoder. For example, in VisualBERT[84], the first N embeddings represent textual information, while the following K embeddings represent visual information separated by a [SEP] token. This approach allows the model to process and interpret the fused multimodal data jointly. Building upon this approach, LXMERT[159] enhances the model’s capacity to extract both intramodal and intermodal features by incorporating self-attention and cross-attention mechanisms. UNITER[24] further improves multimodal interaction learning by combining four training objectives: masked language modeling (MLM), masked region modeling (MRM), word-region alignment (WRA), and image-text matching (ITM).
@@
In the field of remote sensing, RSGPT[55] integrates visual and language modalities using a Q-Former to fuse modality-specific embeddings. It introduces image-question-answer triplets based on remote sensing imagery interpretation as domain-specific knowledge and proposes RSIEval as a benchmark for evaluating the performance of remote sensing multimodal large language models (RSMLLMs). Building on this work, GeoChat[74] fine-tunes a pretrained LLaMA[161] model using four types of visual-language understanding tasks. It distinguishes visual and language modalities using <im_start> and <im_end> tokens, significantly improving the model’s understanding of remote sensing imagery. Similarly, RS-LLaVA[13] adopts the multimodal architecture of LLaVA[94], introducing a projection layer to map visual modalities. Remote sensing images and their descriptions are encoded as image tokens and text tokens, respectively, and the model is further optimized using LoRA[52] fine-tuning techniques.
@@
As shown in Fig.7(b), dynamic hybrid embedding refers to the approach in which data from different modalities are concatenated and directly input into the embedding layer for joint processing. Unlike the modality-tagging method, in dynamic hybrid embedding, the number and positions of the embeddings from each modality are dynamically variable. This dynamic approach allows the model to effectively capture interaction patterns across modalities, thereby enhancing its understanding of multimodal joint representations. For example, OSCAR[87] directly combines textual data, object labels, and object images into a single input, which is uniformly processed through the embedding layer. This approach significantly improves the model’s ability to extract target information in query-based tasks. PaLM-E[35] builds on this work by integrating continuous data streams from the physical world, enabling the model to interact with real-world environments. VIMA[64] focuses on the domain of embodied intelligence and directly combines text tokens, object tokens, and action tokens, allowing the model to learn how to generate appropriate action feedback in response to different instructions and goals, thus enabling robots to perform a wide variety of tasks.
@@
In the field of remote sensing, TEOChat[59] employs a shared vision-language connector to achieve hybrid embedding of image tokens and text tokens, thereby further enhancing the fusion of multimodal information. To improve the model’s understanding and reasoning ability with respect to geographic knowledge, SkySenseGPT[108] incorporates more detailed instructions into text tokens and introduces both region-level and image-level scene graph information. These features significantly strengthen the model’s ability to understand both local and global relationships among geo-objects. Through this approach, remote sensing models are better equipped to capture spatial information and object relationships across different scales and levels, thereby improving their overall ability to intelligently interpret remote sensing imagery.
@@
The shift from modality-tagging methods to joint embedding strategies has become a major development trend for improving the performance of language-centered multimodal models. The core idea is to increase the model’s learning difficulty by disrupting the order of tokens from different modalities, thereby encouraging a deeper understanding of multimodal data. Modality-tagging methods assign fixed positions and lengths to different modalities, allowing the model to clearly distinguish the source of each token. However, this structured design can limit the model’s ability to capture complex relationships across modalities. In contrast, joint embedding mixes and concatenates tokens from different modalities in a less constrained manner, compelling the model to learn intermodal interactions and complementary information within a more flexible and intricate input structure. As a result, the model must not only understand the individual characteristics of each modality but also learn how to integrate latent relationships across modalities, thus enhancing its overall ability to interpret multimodal data holistically.
@@ III-B2 Semantic Alignment across Modalities
Aligning representations from different modalities at the semantic level is a fundamental challenge in multimodal alignment research. Existing approaches to semantic alignment can be broadly categorized into two types: explicit alignment and implicit alignment. The key difference between them lies in whether the alignment relationships across modalities are explicitly introduced during training, such as through additional manual annotations or predefined alignment schemes.
@@
To effectively embed different modalities into the language modality, a straightforward approach is to use paired data–such as image-text or point cloud-text datasets–to train the model. Within this training framework, the model learns to align modalities by modeling the pairing relationships between multimodal data and corresponding textual descriptions. Based on the granularity of alignment between nonlinguistic and linguistic modalities, current research efforts can be broadly classified into three main approaches: scene-caption alignment, region-phrase alignment, and scene-instruction alignment. The characteristics of each approach are illustrated in Fig.8.
@@
  • Scene-Caption Alignment: Scene-caption alignment methods focus on aligning an entire image with its corresponding textual description, aiming to help the model learn the overall association between visual content and language, as illustrated in Fig.8(a). A representative work in this category is CLIP[134], which leverages contrastive learning to achieve effective alignment and has inspired a series of follow-up methods, including UniCL[80], K-LITE[144], and LaCLIP[36]. Owing to the task-agnostic nature of contrastive learning, these approaches are well suited as foundational models that can be quickly adapted to various downstream tasks via fine-tuning, few-shot, or even zero-shot learning. However, these methods face several limitations. First, they rely heavily on large-scale, high-quality image-text paired datasets, and their image understanding ability tends to be shallow. Second, such methods often struggle to handle both image understanding and text generation tasks simultaneously.
    To address these issues, SimVLM[174] introduced the PrefixLM pretraining objective, which maintains strong image understanding while enabling text generation capabilities. Similarly, BLIP[81] performs well in terms of both image understanding and captioning, and its unique CapFilt mechanism effectively reduces noise in the training data. Despite these improvements, scene-caption-based methods generally lack the ability to deeply and finely understand image content. This limitation largely stems from the fact that simple image-text pairs fail to provide sufficiently rich guidance, thereby constraining the model’s capacity to learn more complex vision-language relationships.
@@
    In remote sensing, vision-language models have rapidly become a research focus because of their superior capabilities in image understanding and language processing. One of the early efforts in scene-caption alignment for remote sensing was conducted by Qu et al.[133], who utilized the UCM-Captions dataset[133]. However, this dataset was relatively small in scale and contained simple descriptions, limiting its ability to represent complex content and diverse scenarios in remote sensing imagery. To address these limitations, subsequent works such as RSICD[106] and RSITMD[192] constructed larger, more diverse, and semantically richer datasets while also developing task-specific models for various downstream applications. Building upon these foundations, RemoteCLIP[92] and GeoCLIP[164] have further advanced the field. These studies introduced large-scale scene-caption datasets for remote sensing containing up to 1.6 million and 5 million images, respectively, and proposed two foundational models tailored specifically to the remote sensing domain. Compared with earlier efforts, these studies adopt transformer-based architectures and demonstrate a deeper understanding of remote sensing image content.
@@
  • Region-Phrase Alignment: To enhance the model’s fine-grained understanding of images, some studies have explored incorporating detailed correspondences between image regions and individual words or phrases, as illustrated in Fig.8(b). Early work in this research direction–such as ViLBERT[102]-typically relied on external object detectors to identify regions of interest (ROIs), using techniques such as RoIAlign[46] to extract features from these regions and align them with corresponding text tokens. However, this approach has several notable limitations. First, its dependence on pretrained external detectors restricts the range of detectable object categories and directly affects model performance. Second, the inclusion of an additional detection module increases model complexity and reduces training efficiency.
    To address these issues, VL-BERT[153] integrates the detection module into the training process, while ViLT[70] eliminates it altogether, instead using a pure transformer architecture to extract image features. These adjustments significantly improve efficiency while maintaining strong performance. Additionally, some researchers argue that relying solely on ROI features may lead to the loss of valuable background contextual information. In response, strategies such as background blurring, masking, and feature fusion have been proposed to preserve as much background information as possible while extracting regional features.
@@
    In remote sensing, region-phrase alignment has emerged as a compelling research direction, particularly within the context of scene-caption tasks. Early studies primarily adopted traditional encoder-decoder architectures and focused on extracting features from target regions for image caption generation. While these approaches proved effective, they often overlooked domain-specific knowledge unique to remote sensing. With the rapid advancement of vision-language models, concept-level region-phrase alignment has seen new breakthroughs in remote sensing. For instance, GeoVG[157] was the first to introduce the visual grounding task into the remote sensing domain. It employs a dual-stream architecture to separately extract visual features and geographic relational graphs, which are then integrated through a fusion module to effectively combine geospatial relationships with visual information. Building upon this work, RSVG[195] introduced the MGVLF module, which enhances grounding accuracy by filtering out noise and reinforcing salient features. Additionally, few-shot object detection (FSOD) has become an emerging topic in remote sensing. A recent work, TEMO[105], has made innovative contributions in this direction by incorporating detailed descriptions of all object classes as additional features, effectively mitigating class confusion.
@@
  • Scene-Instruction Alignment: While the previously discussed scene-caption and region-phrase alignment methods effectively achieve direct alignment between visual and language modalities, they often fall short in exploring the complex and abstract relationships among concepts expressed across modalities–an aspect that is crucial for achieving higher-level compositional generalization. Against this backdrop, scene-instruction-based alignment demonstrates its unique value. Its core lies in answering questions through logical reasoning, as illustrated in Fig.8(c). This approach requires the model not only to understand individual concepts but also to grasp the intricate relationships between them.
    In recent years, LLMs have demonstrated powerful general knowledge understanding and reasoning capabilities, which has driven a new research direction that leverages LLMs to address multimodal tasks. By fine-tuning LLMs with scene-instruction datasets, as demonstrated in LLaVA[94] and InstructBLIP[31], models can further enhance their logical reasoning and instruction-following abilities. These works have also simplified the process of constructing high-quality instruction datasets. Moreover, recent studies have introduced novel instruction processing mechanisms; for example, CAT[169] expands narrow language instructions into vision-language instructions, while SEAL[178] incorporates a visual search mechanism to help the model focus more precisely on key regions within an image.
@@
    In remote sensing, scene-instruction alignment is gradually emerging as a critical technique. It enables nonexpert users to interact with remote sensing imagery using natural language. In RSVQA[99], researchers used OpenStreetMap data and manual annotation to create the first large-scale remote sensing visual question answering (VQA) dataset and proposed a CNN-RNN-based VQA model. Subsequent work advanced the field by developing VQA-TextRS[4], a dataset supporting open-ended question answering, alongside a Transformer-based VQA model. In the same year, Prompt-RSVQA[20] proposed a method for mapping images into the language space and feeding them, along with questions, into BERT[72] to generate answers. These studies indicate that remote sensing research is rapidly converging in the direction of vision-language models, laying the groundwork for future innovations in human-AI interactions and multimodal scene understanding.
@@
In summary, the progression from scene-caption to region-phrase and, more recently, to scene-instruction alignment reflects a gradual deepening of the semantic understanding of geo-objects within remote sensing intelligent interpretation. This evolution corresponds to a shift from coarse to fine grained image-text matching strategies. Specifically, scene-caption alignment focuses on matching entire images with holistic textual descriptions; region-phrase alignment refines the image side by dividing it into multiple regions and aligning each with a corresponding phrase; scene-instruction alignment further enhances the granularity of image-text alignment by using complex instructions to describe the relationships among geo-objects. The development of image-text alignment methods from coarse to fine granularity has accelerated the evolution of remote sensing interpretation models from shallow semantic understanding toward deep mechanistic comprehension.
@@
While explicit alignment methods have achieved remarkable success in multimodal learning, they also present several practical limitations. First, they rely on large-scale paired image-text datasets for model training, which poses significant challenges in terms of data collection and preprocessing. Second, achieving alignment between two different modalities typically requires separate training procedures for each modality, increasing both computational and time costs. More importantly, as the number of modalities increases, even when a language-centered alignment strategy is used, the alignment process becomes exceedingly complex and cumbersome. Moreover, when interaction is required between two nonlanguage modalities, using language as an intermediary can lead to issues in both efficiency and accuracy. If the alignment between one modality and language is imperfect, errors may accumulate during the conceptual transfer process, ultimately leading to suboptimal or incorrect interaction outcomes.
@@
Despite the significant heterogeneity across different modalities, they ultimately describe the same physical world. This fundamental fact implies that, regardless of whether the alignment between visual and language modalities is explicitly defined, an intrinsic condition for unified convergence should exist. In support of this condition, Minyoung Huh et al.[57] experimentally demonstrated that visual models with different architectures and training objectives exhibit a high degree of representational consistency and that this consistency continues to improve as the model size and training data scale increase. This pattern also holds for cross-modal vision-language models. Similarly, Maniparambil[113] reported that it is possible to perform subtitle matching or retrieval tasks using unaligned visual and language encoders, thereby further validating the existence of an inherent consistency between visual and language representation spaces. These findings raise a fundamental question: Is there an implicit, inherently defined alignment between language and vision? If so, then it may be possible to achieve concept-level alignment between vision and language even when training on single-modal data—this is the central idea behind implicit alignment, which has recently garnered increasing research interest.
Notably, many of these studies choose to align visual representations to the language space rather than the other way around. This observation leads to a second key question: Why is language often chosen as the convergence point for vision-language alignment? This paper argues that, as the primary medium of human communication, language most accurately reflects human information cognition and logical reasoning mechanisms. Furthermore, language is inherently abstract, generalizable, and adaptable, making it well suited as a global workspace for efficiently organizing a conceptual space that encompasses visual and other nonlinguistic modalities. Building on this view, this paper refers to such research efforts as language-centered implicit multimodal alignment.
@@
Language-centered implicit multimodal alignment aims to explore how nonlinguistic modality information can be directly aligned with the language representation space in the absence of prior modality-language alignment pairs. In this line of research, the work LM4VisualEncoding[126] revealed that even when trained solely on textual data, the encoder of a large language model demonstrates surprisingly strong performance on both pure visual tasks and multimodal tasks. This finding offers new insights into the use of language models for visual tasks, highlighting their ability to encode and process visual-language information through shared parameters. Furthermore, a recent LAMP[1] study confirms the potential of pretrained language models in handling 3D point cloud data, providing additional evidence of the broad adaptability of language models in interpreting and processing nontraditional textual data. Collectively, these research efforts have pushed the boundaries of applying large language models to computer vision tasks and have expanded our understanding of their latent capabilities in processing unimodal visual data.
@@
With the rapid advancement of remote sensing and related technologies, vast amounts of remote sensing data are being generated daily. However, constructing large-scale image-text paired datasets in the remote sensing domain remains a significant challenge. Therefore, introducing implicit alignment into remote sensing is both necessary and timely. Recently, several studies have begun to explore this direction. For example, GRAFT[111] leverages internet images captured from the same ground locations as remote sensing images to serve as an intermediary between remote sensing imagery and language. It proposes a method to train a remote sensing vision-language model without relying on any textual annotations. Such attempts to introduce implicit alignment into the remote sensing field provide promising new directions for addressing the shortage of large-scale multimodal image-text pairs. Nonetheless, this line of research is still in its early stages and requires further exploration.
III-C Language-Centered Knowledge Association and Reasoning